OpenStudy (zoeeh127):

A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks. Number of pages 50 100 150 200 Cost 1.75 2.55 3.35 4.15 Which equation gives the cost, y, in terms of the number of pages, x? A. y = 0.016x + 0.95 B. y = 0.016x + 1.75 C. y = 0.016x + 48.25 D. y = 52.6x + 0.95

OpenStudy (zoeeh127):

it comes out to be 0.80 over 50

OpenStudy (zoeeh127):

but the graph says 1.75 over 50 so i'm lost

OpenStudy (texaschic101):

what is .80/50 in decimal form ?

OpenStudy (zoeeh127):

0.035 for the graph and 0.0016 for the answer i came up with

OpenStudy (texaschic101):

0.80/50 = .016

OpenStudy (zoeeh127):

sorry i added a zero on accident

OpenStudy (zoeeh127):

i'm so confused

OpenStudy (texaschic101):

now use y = mx + b slope(m) = .016 and use either of your points....(50,1.75)...x = 50 and y = 1.75 now sub and find b, the y intercept

OpenStudy (texaschic101):

y = mx + b 1.75 = .016(50) + b 1.75 = 0.8 + b 1.75 - 0.8 = b 0.95 = b so your equation is : y = .016x + .95 understand ??

OpenStudy (texaschic101):

do you have any questions about what I just did ?

OpenStudy (texaschic101):

first, you take 2 sets of points and the slope formula to find the slope. Then you take y = mx + b form and one set of points....sub in your info and you will find the y int...thus, your equation

OpenStudy (zoeeh127):

THANK YOU SO MUCH THAT REALLY HELPED